Total Cost Concept of Product Pricing

Voice Com, Inc., produces and sells cellular phones. The costs of producing and selling 6,000 units of cellular phones are as follows:Total Cost Concept of Product Pricing Voice Com, Inc., produces and sells cellular phones. The costs of producing and selling
Answer #1

Part 1

Total cost

$1320000

Cost amount per unit

$220

Total costs:

Variable costs

984000

Fixed costs (248600+87400)

336000

Total costs

$1320000

Number of units produced

60000

Cost per unit

$220

Variable costs = total variable cost per unit * number of units produced = 164*6000 = 984000

Fixed costs = fixed factory overhead + fixed selling and administrative expenses = 248600+87400 = 336000

Cost per unit = total costs / number of units produced = 1320000/6000 = 220

Part B

Markup percentage

7.80%

Markup percentage = desired profit / total costs

Desired profit = invested assets * rate of return = 735430*14% = 102960.20

Markup percentage = 102960.20/1320000 = 7.80%

Part C

Selling price

$237.16 per phone

Cost per unit

$220

Markup (220*7.80%)

17.16

Selling price

$237.16